# Heterogeneity in Entity Matching: A Survey and Experimental Analysis

### v1

@@ 3.3 Deep Learning and Graph Neural Networks
Seq2SeqMatcher [51] treats EM as a token-level sequence-to-sequence problem. Its align-compare-aggregate neural architecture handles schema and format heterogeneity by resolving attribute mismatches and managing noisy data.
Graph-based models have shown particular promise for heterogeneous settings. LinKG [91] offers a scalable framework for linking heterogeneous entity graphs. It combines LSTM-based encoding for textual data, locality-sensitive hashing for scalability, and heterogeneous graph attention networks to resolve ambiguous links. Its deployment in Microsoft Academic Search highlights its practical effectiveness.
@@
GTA [17] combines graph contrastive learning with Transformers. It constructs hybrid graphs for dual-level matching and multi-granularity interaction, achieving high accuracy by jointly leveraging semantic embeddings and relational structure.
3.4 Leveraging Knowledge Bases
Knowledge bases (KBs) play a critical role in addressing semantic and structural heterogeneity in EM by providing external context and structured representations of entities, relationships, and concepts. Structured KBs such as DBpedia, Wikidata, and domain-specific ontologies help resolve terminology mismatches, schema inconsistencies, and ambiguities in meaning. These sources encode synonymy, polysemy, and hierarchical relationships, enabling more precise alignment of records across heterogeneous datasets. When incorporated into EM pipelines, KBs can support both attribute-level and entity-level matching, enhancing robustness in high-variability settings. Additionally, unstructured or semi-structured sources—such as textual corpora and web data—complement structured KBs by offering further semantic signals useful for disambiguation and contextual matching.
Most work in this space focuses on applying EM techniques within or between knowledge graphs, with less attention given to how external KBs can broadly support EM tasks. Below, we summarize several notable contributions that integrate KBs with EM systems to handle diverse forms of heterogeneity.
CrossKG [75] tackles the problem of entity alignment across different knowledge graphs. It introduces an embedding-based model that maps entities into a shared latent space by generating attribute character embeddings. The method uses transitivity-based enrichment over attribute triples to improve semantic alignment. Experimental results show significant gains in cross-graph alignment accuracy, illustrating its effectiveness in resolving semantic inconsistencies and supporting integration across heterogeneous knowledge graphs.
Ontological Graph Keys (OGKs[43] extend conventional graph key techniques to support ontological subgraph matching. Unlike syntactic key matching, OGKs leverage external ontologies to detect semantically equivalent but syntactically divergent entities. The authors also introduce scalable algorithms, including a budgeted Chase procedure to efficiently manage computation costs. This approach addresses semantic heterogeneity and attribute-level inconsistencies, making it well-suited for complex EM scenarios.
Bornemann et al. [7] examine a more nuanced form of semantic heterogeneity by exploring role alignment in temporal knowledge bases. Their framework models entities with dynamic roles—e.g., “Joe Biden” as both President and Commander-in-Chief—using time-stamped constraints to infer equivalence over time. This allows EM systems to incorporate temporal reasoning when resolving evolving entity states.
CollectiveEA [90] proposes a collective entity alignment framework that incorporates structural, semantic, and string-based similarity features. It treats entity alignment as a stable matching problem and solves it using the deferred acceptance algorithm. The model is tested on both cross-lingual and monolingual benchmarks, showing strong results and demonstrating its ability to address both structural and semantic heterogeneity in knowledge-graph-based EM.
@@ 3.5 Transfer Learning and Domain Adaptation
Transfer learning and domain adaptation have become powerful tools for tackling HEM, particularly in scenarios with limited labeled data or significant domain shifts. Transfer learning enables models to reuse knowledge learned from large, general-purpose datasets by fine-tuning them on smaller, domain-specific EM tasks. This helps in aligning terminology, language, and semantic patterns across datasets. Domain adaptation complements this by explicitly addressing discrepancies between source and target domains—such as differences in data schemas, vocabulary, or structure—using techniques like adversarial training, domain-specific embeddings, and distribution alignment. Together, these methods reduce the need for task-specific supervision while improving generalization across heterogeneous sources.
Auto-EM [93] introduces a transfer learning framework for EM by leveraging deep models pre-trained on large-scale knowledge bases. It uses a hierarchical neural network to pre-train entity-specific models (e.g., for locations, organizations) using rich synonyms and contextual information. These models can be fine-tuned or directly applied to new tasks, reducing reliance on labeled data. Auto-EM effectively aligns semantically similar entities and adapts across entity types, demonstrating strong performance on diverse EM benchmarks.
AdaMEL [33] presents a deep transfer learning method for multi-source entity linkage. It employs attribute-level self-attention to capture the importance of individual features, while domain adaptation mechanisms allow generalization across different distributions. The model supports integrating labeled data from multiple domains, enhancing accuracy and robustness. AdaMEL effectively addresses both semantic heterogeneity (e.g., terminology shifts) and representation heterogeneity (e.g., format or attribute variation).
@@ 3.6 Active Learning and Interactive Methods
Active learning is particularly relevant to HEM due to the increased need for labeled data in complex, heterogeneous environments. Unlike standard EM settings where modest supervision often suffices, HEM demands targeted labeling to handle diversity in schemas, formats, and semantics. Active learning addresses this by selecting the most informative or uncertain record pairs—using strategies like uncertainty sampling or committee-based selection—to maximize model performance with minimal annotation cost.
@@
DIAL [32] proposes a scalable active learning strategy using an Index-By-Committee framework. It jointly optimizes blocking recall and matching precision, using pre-trained transformers to compute semantic embeddings. DIAL scales to large Cartesian product spaces and shows strong results on multilingual datasets, tackling both semantic and format heterogeneity.
[23] frame EM as a progressive labeling process, where an oracle (human or system) labels record pairs on demand. Their algorithms optimize recall under a fixed query budget, creating an adaptive feedback loop. Extending this, [68] develop a system for querying EM results on specific data subsets without executing a full pipeline. The system dynamically constructs and updates indices to enable low-latency, user-driven exploration.
@@
CollaborER [26] introduces a self-supervised framework that generates pseudo-labels and trains matchers collaboratively. It integrates graph and sentence-level features, outperforming existing unsupervised baselines and rivaling supervised models—while addressing both semantic and structural heterogeneity.
ALWUS [8] focuses on unstructured datasets, applying uncertainty sampling to select training pairs. It incorporates NLP features like soft TF-IDF and supports various classifiers, including SVMs and deep networks. ALWUS yields strong performance with limited supervision, particularly on noisy, text-heavy datasets.
Finally, JedAI 2.0 [57] offers an end-to-end, interactive entity resolution platform. Its GUI allows users to configure workflows, visualize matches, and refine matching strategies. JedAI supports schema-free and loosely structured data, improving usability and performance for HEM in practical scenarios.
3.7 LLMs

### v2

@@ 3.3 Deep Learning and Graph Neural Networks
Seq2SeqMatcher [85] treats EM as a token-level sequence-to-sequence problem. Its align–compare–aggregate architecture handles schema and format heterogeneity by resolving attribute mismatches and managing noisy data.
Graph-based models have shown particular promise in heterogeneous settings. LinKG [146] offers a scalable framework for linking heterogeneous entity graphs. It combines LSTM-based encoding for textual data, locality-sensitive hashing for scalability, and heterogeneous graph attention networks to resolve ambiguous links. Its deployment in Microsoft Academic Search highlights its practical effectiveness.
@@
GTA [29] integrates graph contrastive learning with Transformers. It constructs hybrid graphs for dual-level matching and multi-granularity interaction, achieving high accuracy by jointly leveraging semantic embeddings and relational structure.
@@
3.4 Knowledge Graphs and Ontologies for EM and Heterogeneous EM
Knowledge Graphs (KGs) and ontologies have long been used in data integration and linkage [72, 11], and they provide structured, semantically rich representations of entities, attributes, and relationships that can support EM under both semantic and structural heterogeneity. By encoding synonymy, polysemy, hierarchical relations, and domain constraints, KGs enable EM systems to reconcile terminology mismatches, disambiguate ambiguous attributes, and exploit contextual signals not explicitly present in raw records. When incorporated into EM pipelines, KGs can support attribute-level and entity-level matching, improving robustness in noisy or heterogeneous environments.
KGs also help mitigate representation heterogeneity by offering schema-independent cues. For example, mappings between concept hierarchies or ontological types allow EM systems to compare records even when their schemas differ or when attribute names are misaligned. Unstructured or semi-structured sources—such as text corpora, enterprise metadata, or domain-specific ontologies—further complement structured KGs by providing latent semantic information that can be integrated through embedding models.
Several contributions illustrate how KGs can directly enhance EM. Ontological Graph Keys (OGKs[72] extend classical graph-key approaches by leveraging external ontologies to detect syntactically divergent but semantically equivalent subgraphs. Their scalable budgeted-Chase–based algorithms allow EM systems to reconcile attribute-level inconsistencies under semantic heterogeneity. Temporal KGs provide another important signal for EM: Bornemann et al. [11] align entities with evolving roles using time-stamped constraints, enabling EM systems to incorporate temporal semantics when matching entities whose descriptions change over time.
Integrating external knowledge resources—including ontologies, domain schemas, and large KGs such as DBpedia and Wikidata—provides powerful semantic context that can improve both the accuracy and explainability of EM, particularly in settings marked by terminology variation, contextual ambiguity, or schema drift.
3.5 Ontology Matching and KG Alignment Under Heterogeneity
Ontology matching and knowledge graph (KG) alignment are well-established areas [53, 52, 114] that address heterogeneity at the schema, concept, and entity levels. Although distinct from EM, these areas confront many of the same challenges—terminology variation, language differences, granularity mismatches, contextual ambiguity, and structural divergence. Techniques developed for ontology and KG alignment therefore offer valuable insights for building heterogeneity-aware EM systems.
Traditional ontology matching methods such as LogMap [53], PARIS [114], and AgreementMakerLight (AML) [52] focus on aligning classes, properties, and schema elements across ontologies using combinations of lexical cues, structural constraints, and logical reasoning. These systems explicitly target representation heterogeneity by reconciling divergent schema structures and terminologies across domains. PARIS, in particular, demonstrates how instance-, schema-, and lexical-level evidence can be integrated through probabilistic reasoning to address multi-level heterogeneity.
Recent KG-alignment methods extend these ideas to entity-level alignment in large-scale, often multilingual knowledge graphs. Early embedding-based models such as MTransE [15] and BootEA [118] learn shared latent spaces that align structurally similar entities across heterogeneous KGs. Subsequent approaches such as RDGCN [138] leverage Graph Neural Networks to integrate structural, textual, and relational contexts, while CrossKG [122] enriches attribute information using attribute triples, character-level embeddings, and transitivity rules. Methods such as RREA [74] and MultiKE [47] further incorporate multi-view or relation-aware representations to capture semantic variation, language differences, and schema divergence.
CollectiveEA [145] combines structural, semantic, and lexical signals to align entities across heterogeneous KGs. It extracts graph-neighborhood features, textual cues, and string similarities, and performs collective alignment through a stable matching formulation that considers interdependencies between entities. CollectiveEA exemplifies how multi-source evidence can be integrated to address substantial structural and semantic heterogeneity in large KGs.
Although ontology and KG alignment target different tasks from EM, they share key goals: reconciling mismatched representations, resolving terminology variation, and establishing semantic equivalence across heterogeneous sources. Techniques from this literature—including cross-lingual embeddings, structure-aware reasoning, and multi-view alignment—offer promising directions for future heterogeneity-aware EM systems.
3.6 Benchmarks for HEM
Benchmarks play an essential role in evaluating EM systems under different forms of heterogeneity [26, 134, 40, 21, 15], yet many widely used datasets focus primarily on clean, schema-aligned tables with limited variation in semantics, granularity, or representation. For HEM, benchmarks must expose EM models to realistic sources of heterogeneity—including terminology variation, schema drift, inconsistent attribute granularity, noisy or missing values, multilingual data, and domain shifts. Several recent benchmark suites and evaluation frameworks explicitly address these challenges.
The Magellan family [26, 128] introduced a set of diverse EM tasks spanning structured, web-derived, and enterprise datasets. Although not originally designed for heterogeneity, many Magellan datasets contain natural representation variability, schema inconsistencies, and string-level noise, making them suitable for evaluating blocking, similarity-based matching, and supervised learners under non-uniform data conditions. Subsequent extensions such as AutoBlock [128] provide large, heterogeneous blocking datasets with realistic attribute misalignments.
WDC Web Table Matching [134] and WDC Product Matching datasets contain highly heterogeneous web-extracted product data with significant noise, inconsistent taxonomies, missing values, and differing attribute vocabularies. These datasets directly test robustness to terminology heterogeneity, representation drift, and unstructured data integration, and are widely used in recent deep EM papers such as DeepMatcher and Ditto.
GEM and GEMBench [40] were specifically designed to evaluate generalization in EM models across domains, schemas, and modalities. The datasets span multiple domains with different schemas, attribute distributions, and linguistic styles, exposing deep EM methods to cross-domain and cross-schema heterogeneity. GEMBench additionally provides systematic splits for in-domain, out-of-domain, and zero-shot evaluation, aligning directly with representation, contextual, and semantic variability.
More recently, LLM-oriented EM benchmarks such as PromptEM and MatchGPT [21] include natural-language–rich attributes, schema inconsistencies, and diverse domains. These datasets are constructed to test reasoning-based matching and robustness to contextual and semantic heterogeneity. They also highlight new sources of error from long free-text attributes, entity descriptions, and domain-shift scenarios.
Finally, multilingual EM datasets—such as the multilingual WDC collections and cross-lingual KG alignment datasets (e.g., DBP15K) [15]—provide natural test beds for evaluating heterogeneity induced by language differences, polysemy, and culturally specific schemas. Although DBP15K predates many recent benchmarks, it remains a standard evaluation dataset for cross-lingual and cross-schema heterogeneity.
The growing ecosystem of heterogeneous EM benchmarks reveals that models must handle not only noise and missingness, but also schema drift, semantic inconsistency, domain shift, and multilingual variation. These benchmarks provide essential test beds for evaluating heterogeneity-aware EM methods and highlight the need for robust generalization beyond narrow, dataset-specific conditions.
3.7 Instance Coreference Resolution and HEM
Instance coreference resolution (CR) is the task of identifying textual or semi-structured mentions that refer to the same real-world entity [49, 112, 64, 55]. Although traditionally studied in natural language processing, CR addresses many of the same heterogeneity challenges that arise in EM: variability in surface forms, contextual ambiguity, differences in granularity, missing or partial descriptions, and cross-lingual variation. As such, CR provides a rich and mature body of techniques that can inform heterogeneity-aware EM.
Classical CR systems [49, 112] rely on string-based, syntactic, and rule-driven cues to determine whether two mentions are coreferent. These systems must resolve substantial representation heterogeneity, as the same entity may be expressed through names, aliases, pronouns, definite descriptions, or abbreviations. Modern neural CR methods significantly extend this capability. The end-to-end neural coreference model of Lee et al. [64] jointly learns mention detection and coreference scoring using contextual embeddings, enabling the model to infer entity equivalence from latent semantic signals rather than explicit surface overlap. SpanBERT-based models [55] further improve robustness by capturing fine-grained semantic similarity across mentions with rich contextualized representations. These neural models show strong ability to reconcile heterogeneity in terminology, context, and syntax—issues that also arise prominently in EM.
Beyond unstructured text, recent work extends CR to semi-structured and multimodal data. For example, resolving product or organization mentions in web tables, listings, XML records, and documents enriched with metadata requires integrating textual cues with structured context, schema information, or visual features [63, 2]. Multimodal CR systems combine text, layout, and image signals to identify cross-mention equivalence in documents, illustrating how heterogeneous data sources can be jointly leveraged when surface similarity is insufficient. Such approaches demonstrate that CR methods increasingly operate in settings where heterogeneity is similar in nature to EM: fragmented or noisy representations, schema variation across sources, differing attribute granularity, and contextual shifts between mentions.
Despite targeting different tasks, CR and EM share core objectives: resolving whether two heterogeneous representations correspond to the same underlying entity. CR’s long-standing emphasis on modeling contextual semantics, handling ambiguity, integrating multiple modalities, and performing global consistency reasoning offers valuable methodological insight for EM under heterogeneity. Advances in CR—particularly in representation learning, contextual modeling, and cross-document reasoning—therefore provide promising directions for designing future HEM systems capable of robust performance across diverse and highly variable data sources.
@@ 3.8 Transfer Learning and Domain Adaptation
Transfer learning and domain adaptation have become powerful tools for tackling HEM [148, 54, 121, 131, 123], particularly in scenarios with limited labeled data or significant domain shifts. Transfer learning enables models to reuse knowledge learned from large, general-purpose datasets by fine-tuning them on smaller, domain-specific EM tasks, helping align terminology, language, and semantic patterns across datasets. Domain adaptation complements this by explicitly addressing discrepancies between source and target domains—such as differences in data schemas, vocabulary, or structure—using techniques like adversarial training, domain-specific embeddings, and distribution alignment. Together, these methods reduce the need for task-specific supervision while improving generalization across heterogeneous sources.
Auto-EM [148] introduces a transfer learning framework for EM by leveraging deep models pre-trained on large-scale knowledge bases. It uses a hierarchical neural network to pre-train entity-specific models (e.g., for locations or organizations) using rich synonyms and contextual information. These models can be fine-tuned or directly applied to new tasks, reducing reliance on labeled data. Auto-EM effectively aligns semantically similar entities and adapts across types, demonstrating strong performance on diverse EM benchmarks.
AdaMEL [54] presents a deep transfer learning method for multi-source entity linkage. It employs attribute-level self-attention to capture the importance of individual features, while domain adaptation mechanisms allow generalization across different distributions. AdaMEL integrates labeled data from multiple domains, enhancing accuracy and robustness. It effectively addresses both semantic heterogeneity (e.g., terminology shifts) and representation heterogeneity (e.g., format or attribute variation).
@@ 3.9 Active Learning and Interactive Methods
Active learning has long been explored in EM and related data integration tasks [99, 51, 36, 109, 50], and it is particularly relevant to HEM due to the increased need for labeled data in complex, heterogeneous environments. Unlike standard EM settings where modest supervision often suffices, HEM demands targeted labeling to handle diversity in schemas, formats, and semantics. Active learning addresses this by selecting the most informative or uncertain record pairs—using strategies like uncertainty sampling or committee-based selection—to maximize model performance with minimal annotation cost.
@@
Early work such as [36] frames EM as a progressive labeling process, where an oracle labels record pairs on demand. Their algorithms optimize recall under a fixed query budget, creating an adaptive feedback loop. Extending this idea, [109] develop a system for querying EM results on specific data subsets without executing a full pipeline, enabling low-latency user-driven exploration through dynamically constructed indices.
JedAI 2.0 [91] offers an end-to-end, interactive entity resolution platform. Its GUI allows users to configure workflows, visualize matches, and refine strategies. JedAI supports schema-free and loosely structured data, improving usability and performance for HEM in practical scenarios.
@@
DIAL [51] proposes a scalable active learning approach using an Index-By-Committee framework. It jointly optimizes blocking recall and matching precision, using pre-trained transformers to compute semantic embeddings. DIAL scales to large Cartesian product spaces and performs well on multilingual datasets, addressing both semantic and format heterogeneity.
CollaborER [41] introduces a self-supervised framework that generates pseudo-labels and trains matchers collaboratively. It integrates graph- and sentence-level features, outperforming existing unsupervised baselines and rivaling supervised models—while addressing both semantic and structural heterogeneity.
@@
3.10 Progressive and Incremental EM and Resolution
Progressive ER methods aim to return high-quality matches early by prioritizing candidate pairs under time or budget constraints. These approaches differ from active learning because they do not rely on human feedback; instead, they incrementally refine similarity scores or ranking functions as additional evidence becomes available. This makes them highly relevant to HEM, where schema drift, missing attributes, datatype inconsistencies, and heterogeneous formats degrade blocking quality and enlarge the candidate space. Classical work such as progressive duplicate detection [92] and large-scale systems like BigDansing [59] demonstrate how adaptive ordering can improve efficiency and robustness in heterogeneous environments.
Subsequent “pay-as-you-go” approaches [75] extend these ideas by adaptively refining similarity signals when schema drift or missing attributes reduce the reliability of individual features. More recent progressive frameworks explicitly model uncertainty arising from heterogeneous or incomplete signals. PERC [57] dynamically reorders comparisons as similarity evidence changes, providing strong performance when attribute-level cues are noisy or inconsistent. Together, these methods align naturally with HEM because representation and semantic heterogeneity make full Cartesian matching impractical and amplify uncertainty in similarity estimation. Progressive ER therefore offers a principled mechanism for coping with heterogeneity-induced ambiguity and computational cost.
3.11 Self-Supervised and Pseudo-Label EM
Self-supervised EM methods learn record representations or matching functions without requiring labeled pairs, making them well suited for heterogeneous settings where supervision is scarce or non-transferable. Heterogeneity across schemas, formats, and vocabularies often limits the applicability of labeled data, and self-supervision offers a way to extract domain-agnostic structure from raw records. EmbDI [14] exemplifies this direction by constructing training corpora from random walks over token–attribute–tuple graphs to capture structural and semantic consistency across heterogeneous tables. More recent approaches such as CollaborER [41] leverage graph- and sentence-level signals to generate pseudo-labels and train matchers collaboratively, reducing reliance on human supervision while addressing both semantic and structural heterogeneity.
Contrastive pretraining methods further strengthen robustness to heterogeneity by learning similarity functions from augmented or multi-view representations. Related contrastive and autoencoder-based approaches (e.g., [39, 95]) learn embeddings resilient to missing values, terminology variation, and schema mismatch. Sudowoodo [132] extends this direction by using contrastive learning to model semantic similarity across noisy, structurally inconsistent textual attributes. These self-supervised strategies directly address core HEM challenges by reducing reliance on labeled data and producing representations that generalize across heterogeneous sources.
3.12 Evolutionary, Meta-Heuristic, and Hybrid Approaches
Evolutionary computation (EC) and meta-heuristic search methods have long been used to optimize complex EM pipelines [135, 149], particularly in settings where heterogeneity renders manually designed rules or fixed classifier parameters ineffective. EC refers to a family of population-based algorithms inspired by biological evolution—including Genetic Algorithms (GA), Genetic Programming (GP), and evolutionary strategies—which explore large combinatorial spaces of matching rules, similarity functions, and blocking configurations. By encoding matching configurations (e.g., attribute selections, similarity metrics, weight vectors) as chromosomes, EC-based methods iteratively evolve high-performing solutions under objectives such as F1 score, precision–recall balance, and coverage, enabling adaptive tuning across heterogeneous schemas, formats, and semantic variations.
Early work by [135] showed how matching rules can evolve over time to adjust to dynamic or heterogeneous environments. Genetic programming approaches such as ERGP [116] evolve composite similarity functions by combining heterogeneous attribute-level comparisons. Beyond individual rules, evolutionary search has been applied to full EM pipelines: [126] optimize attribute-weight configurations in ontology-based product matching to account for schema and terminology drift, and [75] use GA optimization to tune blocking strategies, thresholds, and similarity metrics in a pay-as-you-go framework.
Recent advances extend EC to multi-objective and hybrid optimization settings. For example, multi-objective evolutionary algorithms have been used to jointly optimize accuracy, interpretability, and computational cost, or to evolve interpretable rule sets that complement deep neural matchers [149]. These hybrid EC–ML systems demonstrate that evolutionary optimization can enhance the robustness, explainability, and resource-awareness of learning-based EM pipelines in heterogeneous data environments.
More broadly, the EM literature has recently seen the rise of hybrid architectures that combine heterogeneous forms of evidence—rules, similarity features, blocking signals, schema information, deep embeddings, or even LLM outputs—into unified EM systems. Examples include hybrid symbolic–neural EM frameworks such as DeepER and DeeperFlow [42], hybrid blocking methods that integrate token-based and embedding-based signals (e.g., DeepBlocker [119]), and systems that blend schema-based alignment with neural models for instance matching (e.g., HERA [83]). Although distinct from evolutionary computation, hybrid EM systems share the goal of combining diverse matching signals and modeling paradigms to overcome the limitations of any single technique, and thus naturally relate to EC approaches within the broader landscape of heterogeneity-aware EM.
In comparison to purely neural EM approaches, evolutionary and hybrid methods exhibit complementary strengths and limitations under different forms of heterogeneity. Evolutionary and meta-heuristic techniques are particularly effective in settings dominated by schema, structural, or granularity heterogeneity, where explicit control over attribute selection, similarity functions, and blocking strategies provides flexibility and interpretability. Hybrid architectures further benefit from combining symbolic rules, schema signals, and learned representations, making them more robust to schema drift, missing attributes, and representation mismatch. In contrast, end-to-end neural approaches typically excel under large-scale semantic and linguistic heterogeneity, especially when sufficient labeled data are available, but may be more sensitive to distribution shift, data quality issues, and changes in schema or record structure. This comparison highlights that evolutionary and hybrid methods are not substitutes for neural models, but complementary tools that are often preferable in heterogeneous, low-supervision, or rapidly evolving data environments.
3.13 Large Language Models